Use the following data: Quantity (Q) 1,000 Ordering Cost (Co)-$50 Holding Cost (Ch) $1 Unit Price (P) $25 (A) Calculate and provide the EOQ. (B) Calculate and provide the total number of orders. (C) Calculate and provide the the reorder cycle period in days. (D) Calculate and provide the total cost. (Note: there is no lead time requirement.) (E) Write a paragraph to explain your findings to management.
Answer #1
A) EOQ = (2*A*O/c)^0.5, where
A = Annual demand in units
O = Ordering cost
c = holding cost per unit in $
EOQ = (2*1000*5/1)^0.5 = 100 Units
B) Total number of orders = 1000/100 = 10 orders
C) Reorder cycle period = 365/10 = 36.5 days
D) Total cost:
Material cost = 1000*25 = 25000
Ordering cost = 10*50 = 500
Holding cost = (100/2)*1 = 50
Total cost (including material cost) 25550
E) The EOQ should be 100 units. It would result in the lowest total
ordering cost and holding cost for a year of $550. The firm
would be placing 10 orders and the reordering cycle will be 36.5
days.